OpenStudy (anonymous):

If a car travelling at vm/s takes X meters to stop, what would the stopping distance be if it was travelling at 3vm/s (assume stopping force is equal)

OpenStudy (anonymous):

I remember doing this in class, but forgot which equation to use

OpenStudy (lgbasallote):

hmm could it be \[V_2 ^2 = V_1 + 2as\] where s is the distance

OpenStudy (lgbasallote):

correction.. \[V_2 ^2 = V_1^2 + 2as\]

OpenStudy (anonymous):

V is velocity? a is acceleration? s is displacement?

OpenStudy (lgbasallote):

yup

OpenStudy (anonymous):

Oh... yeah so that's the same one I have here as v^2=u^2+2as

OpenStudy (lgbasallote):

yep same shiz

OpenStudy (lgbasallote):

so first find the acceleration using v = 0 u = v s = X

OpenStudy (anonymous):

0=v^2+2aX... what is a?

OpenStudy (lgbasallote):

you solve for it

OpenStudy (lgbasallote):

isolate a

OpenStudy (anonymous):

a=-v^2/2X

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

now use v^2 = u^2 + 2as again this time use: v = 0 u = 3v a = -v^2/2X

OpenStudy (lgbasallote):

then solve for s

OpenStudy (anonymous):

starting with 0=9v^2+2aX -9v^2/2a=X -9v^2/2(-v^2/2X)=X 9X=X now what?

OpenStudy (lgbasallote):

no...

OpenStudy (lgbasallote):

it's not 2aX

OpenStudy (lgbasallote):

it's 2as

OpenStudy (lgbasallote):

X is just for the first situation

OpenStudy (anonymous):

oh i see

OpenStudy (lgbasallote):

\[0 = 9v^2 + 2as\] \[-9v^2 = 2as\] \[-\frac{9v^2}{2a} = s\] \[-\frac{9v^2}{2(-\frac{v^2}{2X})} = s\] \[\implies -\frac{9v^2}{-\frac{v^2}{X}} = s\] \[\implies 9X= s\]

OpenStudy (anonymous):

Oh. THank You!

OpenStudy (lgbasallote):

welcome